OpenStudy (anonymous):

I need help on finding the answer for the square root of 20y^8.

OpenStudy (anonymous):

2y^4sqrt5

OpenStudy (anonymous):

\[\sqrt{20y^8} = \sqrt{\underline{\color{red}{2 \times 2}} \times 5 \times \underline{\color{red}{y^4 \times y^4}}} = \color{red}2 \sqrt{5} \cdot \color{red}{y^4}\]